Given W(x)=3(2x-4),what is w(5)?
Answer by ikleyn(53937)

.

To answer this question, substitute x= 5 into the formula defining W(x)

    W(5) = 3*(2*5-4) = 3*(10-4) = 3*6 = 18.    ANSWER

It is all they want from you.